Mrs. Jones invested 10,000, some in each of two separate accounts. one pays 5% interest and the other 6%.if her combined annual interest is $575, how much does she have invested in each account?
Answer by mananth(12270)

let $x be invested at 5%
$y be invested @6%
..
interest = $575
..
x+y = 10000.....................1
..
The percentage value is not taken. however the interest is multiplied by 100.
the result is the same
5x+6y=575*100
5x+6y=57,500....................2
multiply equation 1 by -5
-5x-5y=-50,000
add this to (2)
y=7500
$7,500 invested @ 6%
Balance $2500 @ 5%
..
m.ananth@hotmail.ca